# BackdoorMBTI: A Backdoor Learning Multimodal Benchmark Tool Kit for Backdoor Defense Evaluation

### v1

@@ 7. Limititions and Future Work
Mutilmodal application support. BackdoorMBTI aims to provide a unified benchmark with the potential for easy extensions to new modalities, while the modalities supported (i.e., image, text, audio) are integrated individually now. While multimodal applications like visual question answering are not yet supported, we are actively working on it and plan to release this feature in the future.
@@
Scale of Noise Factors. Currently, we only consider Gaussian noise for image and audio data and make simple modifications for text data. We plan to explore additional noise factors in future work.
@@ 9. Acknowledgments
Appendix A Running Experiments
Our experiments are conducted on the GPU server with Intel(R) Xeon(R) Silver 4310 CPU @ 2.10GHz, RTX3090 GPU (24GB) and 128 GB RAM. The versions of all software and packages involved are clearly outlined in the requirements.txt file in the BackdoorMBTI code directory.
Appendix B Hyper-Parameter Settings
Details of the hyper-parameter settings used in our evaluations of backdoor attack and defense algorithms are provided in the framework configuration directory. These settings allow for reproducibility of the reported results.
Appendix C Additional Information of Implemented Backdoor Attacks
Here we introduce the implemented backdoor attacks(BadNets (Gu et al., 2017), BPP (Wang et al., 2022c), SSBA (Li et al., 2021a), WaNet (Nguyen and Tran, 2021), AddSent (Dai et al., 2019), SYNBKD (Qi et al., 2021), LWP (Li et al., 2021d), Blend (Chen et al., 2017), DABA (Liu et al., 2022), GIS (Koffas et al., 2023) and UltraSonic (Koffas et al., 2022)), illustrate the description about what it is and provide necessary information about its implementation.
BadNets: BadNets is the earliest and most renowned backdoor attack method for neural networks. It uses straightforward fixed triggers for all inputs, such as replacing the bottom right corner with a 3x3 white pixel block in images. For text, four uncommon character pairs are randomly selected as triggers and then inserted randomly into the text.
BPP: BPP is a state-of-the-art and sophisticated backdoor attack method targeting machine learning models. It involves injecting tainted data into the training set to manipulate model behavior. This attack is stealthy and exploits the human visual system’s insensitivity to color depth through perturbation techniques such as image quantization and dithering. These perturbations serve as triggers, activating the backdoor under specific conditions.
SSBA: SSBA, short for Sample-Specific Backdoor Attack, stands as a groundbreaking innovation within the domain of poison-only attacks. It marks the first instance of a poison-only sample-specific attack, distinguished by its capability to produce diverse triggers to specific sample inputs through an autoencoder. This refined approach significantly enhances the potency of attack and makes it much more challenging to detect. Furthermore, SSBA’s inherent invisibility and additive attributes amplify its efficacy, enabling seamless integration with legitimate data and evading detection mechanisms with its elevated stealthiness.
WaNet: WaNet is an invisible backdoor attack. Originally conceived as a training control method employing warping-based triggers, WaNet has transformed into a poisoning-based backdoor attack through simple modifications and integrated into the backdoor training pipeline. The training control component has been eliminated, yet the attack continues to yield favorable results.
AddSent: AddSent represents a sentence-level text backdoor attack similar to BadNets, distinguished by its trigger, which is a phrase: ”I watch this 3D movie.” In contrast to BadNets’ trigger based on uncommonly used characters like ’cf’, AddSent employs a complete sentence as its trigger, which is potentially related in meaning to the original context. This semantic alignment contributes to its stealthiness at the meaning level.
SYNBKD: SYNBKD is a sentence-level text backdoor attack that uses text style transfer for adversarial and backdoor purposes, to alter the syntax of a sentence while maintaining its semantic meaning.
LWP: LWP refers to a layer-weight poison backdoor attack, employing straightforward composite triggers instead of pre-defined rare tokens. This approach makes detection challenging within the model’s vocabulary.
Blend: Blend is originally an image attack that utilizes a blending trigger to achieve stealthiness. It has been extended to audio by incorporating a fixed random trigger into the original audio to generate poisoned audio data
DABA: DABA, also known as Opportunistic Backdoor Attack, employs ambient environmental noise as its backdoor trigger mechanism. This attack’s inherent stealthiness, exploiting human auditory habits, often evades detection by systems and users. With deterministic trigger selection, performance-independent sample blending, and trigger-enhanced pipelines, DABA facilitates robust model poisoning, resulting in increased attack success rates.
GIS: GIS, a sophisticated audio backdoor attack focused on style, explores the use of stylistic backdoor triggers and introduces a method for generating dynamic triggers using guitar effects.
UltraSonic: By employing ultrasonic audio signals sampled at 44.1kHz as triggers, this approach achieves imperceptible backdoor injection, imperceivable to the human ear. Due to its fixed sampling rate, we use the same fixed sample rate in our benchmark.
Appendix D Additional Information of Implemented Backdoor Defenses
Here we introduce the implemented backdoor attacks(AC (Chen et al., 2018), STRIP (Gao et al., 2019), FT (Liu et al., 2018a), FP (Liu et al., 2018a), ABL (Li et al., 2021b), CLP (Zheng et al., 2022a) and NC (Wang et al., 2019)), illustrate the description about what it is and provide necessary information about its implementation.
AC: Activation Clustering is based on the observation that clean samples and poisoned samples display distinct activation in the final hidden layer. It identifies backdoor sample clusters through hidden layer activation, facilitating backdoor detection. To identify malicious clusters, various methods have been proposed, such as smaller size, relative size, distance, and silhouette scores. In our benchmark, we use the default criterion of smaller clusters to identify backdoor sample clusters, grounded on the observation that poisoned samples constitute only a small portion of the original dataset.
STRIP: STRIP introduces substantial perturbations into each input and utilizes entropy to identify poisoned inputs. This method relies on the robust characteristic that inputs containing a backdoor attack trigger lead to targeted predictions by attackers. Its universality allows for easy extension to the text and audio domains, making it a widely adopted technique.
FT: Fine-tuning involves using a sanitized subset of the poisoned training dataset to retrain the backdoor model, resulting in a sanitized model. Unlike other methods that solely fine-tune the last layer, we employ fine-tuning across all layers to mitigate the backdoor effect.
FP: Fine-pruning employs neuron pruning to mitigate the impact of backdoors. This method operates under the assumption that clean data and backdoor data follow distinct triggering paths. It prunes inactive malicious neurons based on clean activation paths to eliminate the backdoor. The pruning criterion for FP is based on the activation values of the Batch Normalization layer in the final network layer, although other activation values could also be used as pruning criteria.
ABL: ABL observes that poison data is more readily learned and often converges earlier in training compared to clean data. It identifies poison samples by discerning differences in loss function values during initial training and then maximizes the loss of these isolated data points to achieve backdoor forgetting.
CLP: CLP operates under the assumption that the backdoor neuron demonstrates higher Channel Lipschitzness. It conducts fine-pruning of the Batch Normalization layer using Channel Lipschitzness to achieve a clean model.
NC: NC uses an optimization searching algorithm to identify a short path within different classes for reverse engineering the backdoor trigger. It then uses the discovered possible trigger to determine whether it’s a backdoor model. Finally, it employs fine-tuning to train the model accordingly for mitigation.

### v2

@@ 7. Limititions and Future Work
Multimodal Application Support. BackdoorMBTI aims to provide a unified benchmark with the potential for seamless extension to new modalities. In particular, BackdoorMBTI currently supports a diverse range of single-modality applications, including image, text, audio, video, and contrastive learning, as well as the multimodal application VQA. Although other multimodal applications, such as audiovisual, are not yet supported, we are actively developing this feature and plan to release it in the future.
@@
Scale of Noise Factors. Currently, we only consider Gaussian noise for image and audio data and limited modifications for text data. We plan to explore additional noise factors in future work.
Scalability and Generalizability. BackdoorMBTI enables seamless integration of new modalities, such as video and audio, and ensures robust performance across diverse datasets and tasks. We are expanding our framework to incorporate additional models and attack types that involve different modalities. To facilitate the integration of new modalities, models, attacks, and defenses, we have open-sourced our framework along with detailed documentation.
@@ 9. Acknowledgments
This work was partly supported by the National Key Research and Development Program funded by the Ministry of Science & Technology of China (No. 2023YFB2704903) and the Natural Science Foundation of Shanghai (23ZR1429600).
Appendix A The Design of Noise Generator
The noise generator could generate different noise for varying data. For text data, the noise generator supports 3 distinct strategies:
  • Character-level noise involves operations such as random character substitutions, insertions, or deletions to simulate typographical errors.
  • Word-level noise introduces synonym replacements, shuffling, or random word insertions to emulate semantic variations or adversarial perturbations.
  • Sentence-level noise applies paraphrasing, sentence reordering, or random sentence insertions/deletions to simulate more complex structural changes.
Appendix B The Impact of Low Accuracy on Backdoor Attacks and Defenses
Backdoors can be learned within just a few epochs for most attacks (as observed in (Li et al., 2021b) and reproduced in (Wu et al., 2022)). This suggests that backdoor features are learned strongly, while the model performs poorly on clean class features.
For backdoor defenses, the situation is more complex than backdoor attacks. The key question is whether poor accuracy negatively impacts the effectiveness of various defense strategies.
Only methods that rely on features derived from the dataset or class are likely to be affected by low accuracy. However, due to the quick learning phenomenon of backdoor, it tends to stand out compared to the noise from less-optimized models.
Most defense methods are not impacted by low accuracy, except for those relying on data/class features, where backdoor features remain detectable compared to model noise. Extreme low accuracy (below 70%) is rare in practice, and typical lower accuracies do not significantly degrade defense performance.
Table 6. The impact of low accuracy on backdoor attack. In this experiment, we use BadNets attack on resnet-18 model training on CIFAR-10 dataset, and the default poison ratio is 0.1.
Epochs Clean Accuracy Attack Success Rate
1 0.4941 0.9996
2 0.5421 0.9998
5 0.6938 0.9997
10 0.7501 1.0000
20 0.7900 1.0000
30 0.8164 0.9996
40 0.8264 0.9997
50 0.8377 1.0000
Table 7. The impact of low accuracy on backdoor defense. In this experiment, we select fine-tune as the defense method, the epochs indicates the fine-tuning epoch. We show the clean accuracy (CAC) and attack success rate (ASR) before and after the fine-tune method.
[…]
10 0.4941 0.7713 0.9996 0.0031
10 0.5421 0.7820 0.9998 0.0194
10 0.6938 0.8019 0.9997 0.0057
10 0.7501 0.8183 1.0000 0.0103
10 0.7900 0.8343 1.0000 0.0078
10 0.8164 0.8398 0.9996 0.0385
10 0.8264 0.8468 0.9997 0.02766
10 0.8377 0.8409 1.0000 0.0175
Appendix C The Impact of Noise Level on Backdoor Attacks and Defenses
The parameter noise level is critical in backdoor learning, as noted in BackdoorBench (Wu et al., 2022), varying the factor would provide deeper insights into the impact on evaluation results. We conducted a series of experiments to analyze how varying noise levels affect the effectiveness of backdoor attacks and the robustness of defenses.
Table 8. The impact of noise level on backdoor attack. We used the same experimental settings as those in Table 6, and the number of training epochs was 10.
Mean Variance Clean Accuracy Attack Success Rate
-0.5 0.1 0.8328 0.9998
-0.5 0.5 0.8248 0.9996
-0.5 1.0 0.8241 0.9997
0.0 0.1 0.8356 0.9990
0.0 0.5 0.8274 0.9997
0.0 1.0 0.8185 1.0000
0.5 0.1 0.8274 0.9997
0.5 0.5 0.8287 0.9996
0.5 1.0 0.8212 0.9998
Table 9. The impact of noise level on backdoor defense. We used the same experimental settings as those in Table 7.
Epochs Mean Variance Clean Accuracy Attack Success Rate
10 -0.5 0.1 0.7990 0.1174
10 -0.5 0.5 0.8008 0.1104
10 -0.5 1.0 0.7972 0.1282
10 0.0 0.1 0.7979 0.0727
10 0.0 0.5 0.8019 0.2118
10 0.0 1.0 0.7998 0.0517
10 0.5 0.1 0.8052 0.0615
10 0.5 0.5 0.8018 0.2092
10 0.5 1.0 0.7962 0.1216
Appendix D The Impact of Poison Ratio on Backdoor Attacks
Table 10. The impact of poison rate on backdoor attack.
Pratio Clean Accuracy Attack Success Rate
0.005 0.7746 0.754
0.01 0.7756 0.8529
0.04 0.768 0.9268
0.08 0.7669 0.9486
0.1 0.7633 0.9574
0.2 0.7572 0.9684
0.3 0.7351 0.9762
0.4 0.7208 0.9811
0.5 0.6984 0.985
The poison ratio is a critical factor in determining the effectiveness of backdoor attacks and the robustness of defenses. Varying this parameter is essential to gain a deeper understanding of its impact on attack success rates and defense effectiveness. By systematically adjusting the poison ratio, we can investigate how the prevalence of poisoned data affects a model’s susceptibility to backdoors.